# Application of Langevin Dynamics to Advance the Quantum Natural Gradient Optimization Algorithm

### v3

@@ Abstract
A Quantum Natural Gradient (QNG) algorithm for optimization of variational quantum circuits has been proposed recently. In this study, we employ the Langevin equation with a QNG stochastic force to demonstrate that its discrete-time solution gives a generalized form of the above-specified algorithm, which we call Momentum-QNG. Similar to other optimization algorithms with the momentum term, such as the Stochastic Gradient Descent with momentum, RMSProp with momentum and Adam, Momentum-QNG is more effective to escape local minima and plateaus in the variational parameter space and, therefore, achieves a better convergence behavior compared to the basic QNG. In this paper we benchmark Momentum-QNG together with basic QNG, Adam and Momentum optimizers and find the optimal values of its hyperparameters. Our open-source code is available at https://github.com/borbysh/Momentum-QNG
@@ 1 Introduction
A computationally efficient method for evaluating analytic gradients on quantum hardware has been recently proposed [6]. Therefore, the application of optimization algorithms from the Stochastic Gradient Descent (SGD) family has become possible. However, the path of steepest descent in the parameter space, guided by the (opposite) gradient vector direction, is usually not optimal, because it depends on the number of variational parameters, which is usually excessive. The same overparametrization problem is present in classical Machine Learning (ML) (see, e.g., [7]). To mitigate it, a Natural Gradient (NG) concept has been proposed [8]. Contrary to vanilla SGD, NG determines the steepest descent direction taking into account the Riemannian metric tensor. In this way, the optimization path becomes invariant under arbitrary reparametrization (see [8]) and, therefore, does not suffer from overparametrization [9].
Inspired by the NG approach, Stokes et al. [10] have recently proposed its generalization to the quantum circuit optimization task, which they called a Quantum Natural Gradient (QNG) optimizer. They considered a parametric family of unitary operators $U_{\theta}\in U(N)$, which are indexed by real parameters $\theta\in\mathbb{R}^{d}$. With a fixed reference unit vector $|0\rangle\in\mathbb{C}^{N}$ and a Hermitian operator $H=H^{\dagger}$ acting on $\mathbb{C}^{N}$, they consider the following optimization problem
@@
where $\psi_{\theta}=U_{\theta}|0\rangle$ and $P_{\psi_{\theta}}\in\mathbb{CP}^{N-1}$ is the associated projector. Note that $\psi_{\theta}$ is normalized, since $U_{\theta}$ is unitary. Any local optimum of the nonconvex objective function $\mathcal{L}(\theta)$ can be found by iterating the following discrete-time dynamical system,
[…]
$\left.+\frac{1}{2\eta}\|\theta-\theta_{n}\|_{g(\theta_{n})}^{2}\right]\,,$ (2)
where $\eta>0$ is a positively defined constant, $g(\theta_{n})$ is the symmetric matrix of the Fubini-Study metric tensor $g_{ij}(\theta)=\operatorname{Re}[G_{ij}(\theta)]$, and the Quantum Geometric Tensor is defined as follows (for further details see [10] and references therein):
$G_{ij}(\theta)=\left\langle\frac{\partial\psi_{\theta}}{\partial\theta^{i}},\frac{\partial\psi_{\theta}}{\partial\theta^{j}}\right\rangle-\left\langle\frac{\partial\psi_{\theta}}{\partial\theta^{i}},\psi_{\theta}\right\rangle\left\langle\psi_{\theta},\frac{\partial\psi_{\theta}}{\partial\theta^{j}}\right\rangle\,.$ (3)
In (1), according to [10], we introduced the following notation:
$\|\theta-\theta_{n}\|_{g(\theta_{n})}^{2}=\langle\theta-\theta_{n},g(\theta_{n})(\theta-\theta_{n})\rangle\enspace.$ (4)
Then, the first-order optimality condition corresponding to (1) is:
$g(\theta_{n})(\theta_{n+1}-\theta_{n})=-\eta\cdot\nabla\mathcal{L}(\theta_{n})\enspace.$ (5)
@@
$\theta_{n+1}-\theta_{n}=-\eta\cdot g^{-1}({\theta_{n}})\nabla\mathcal{L}(\theta_{n})\enspace.$ (6)
In this way, with the Fubini-Study metric tensor introduced, the implied descent direction in the parameter space, given by the right-hand side of (6), becomes invariant with respect to arbitrary reparametrization and, therefore, to the details of QNN architecture. With their QNG optimizer (6), the authors of [10] achieved a considerable optimization performance improvement as compared with SGD and Adam [11].
However, being rather effective for convex optimization, QNG often sticks to local minima, saddles and plateaus of non-convex loss functions. In classical ML applications, employing optimization algorithms with momentum (inertial) term, such as SGD with momentum [12], RMSProp with momentum [13] and Adam [11], has demonstrated better convergence characteristics.
Recently, Borysenko and Byshkin [14] demonstrated that a discrete-time solution of the Langevin equation with stochastic gradient force term results into the well-known SGD with momentum [12] optimization algorithm. In this paper, we consider the Langevin equation with the QNG stochastic force term. In Section 2 we show that its discrete-time solution yields a generalized QNG optimization algorithm, which we call Momentum-QNG. In Section 3 we benchmark it together with the basic QNG, Momentum and Adam on several optimization tasks to demonstrate its enhanced performance.
2 Langevin equation with QNG force
The adaptation of Langevin dynanimcs for optimization suggests a new prospective research direction (see, e.g., [15] and [16] ). For optimization of analitically defined objective functions, even a quantum form of the Langevin dynamics has been proposed [17]. In Langevin dynamics, two forces are added to the classical Newton equation of motion – a viscous friction force proportional to the velocity with a friction coefficient $\gamma\geq 0$ and a thermal white noise. Explicitly, the Langevin dynamics of a (Brownian) particle with unit mass $m=1$ in the space of real variables $\theta\in\mathbb{R}^{d}$ and real time $t$ can be described by the following equation (see e.g. [18, 19, 20, 21]):
@@ 2 Langevin equation with QNG force
where $v=d\theta/dt$ denotes velocities and $R(t)$ is a random uncorrelated force with zero mean.
The discrete-time form of (7) with stochastic force $\hat{f}=f+R$ reads:
@@
from (9) one obtains a generalized form of (6):
@@
which we call Momentum-QNG. Note, that (13) reduces to (6) for $\rho=0$.
From (13) one can see that our Momentum-QNG optimization algorithm is a quantum adaptation of SGD with momentum [12] and should not be confused with the recently introduced Momentum QNG algorithm [22] (which is a quantum adaptation of Adam [11]) proposed by the authors of qBang [22].
3 Results and discussion
In this section we benchmark Momentum-QNG with reference to other gradient-based algorithms, such as basic QNG, Adam and Momentum optimizers. In all calculations we set $\Delta\theta_{0}=0$ as the initial condition.
@@ 3.1 Variational Quantum Eigensolvers
Our first test-drive model is an Investment Portfolio Optimization task, which can be mapped to the $N$-particle ($N$ is a number of companies in portfolio) Ising spin-glass model. The ground state of the corresponding $N$-qubit Hamiltonian is parametrized with the VQE ansatz and further optimized to find the energy minimum. Here we explore the cases $N=6$ and 12. To compare the performance of different optimizers, we run a series of 200 trials on a modified tutorial code by Chi-Chun Chen [23]. Each trial is initialized with the random-guess values of variational parameters, being the same for all optimizers. Next, the optimization process runs for 200 steps or until energy convergence up to the 3-digit accuracy. The values of hyperparameters are as follows: $0.01\leq\eta\leq 0.25$, $\rho=0.9$ for Momentum and Momentum-QNG, and $\beta_{1}=0.9$, $\beta_{2}=0.99$, $\epsilon=10^{-8}$ for Adam. To compare the performance of different optimizers, we calculate the mean and standard deviation values of $\Delta E=E_{\text{opt}}-E_{\text{ground}}$ – the difference between the optimized and the exact ground state energy. From Fig. 1(a) one can see that for $N=6$ Momentum-QNG gives the least mean $\Delta E$ between all the optimizers studied over the whole range of stepsize (learning rate $\eta$) values under consideration. On Fig. 1(c) for $N=12$ the situation is similar, except that for $\eta>0.1$ Momentum-QNG tends to divergence. At the same time, the least mean $\Delta E$ achieved by Momentum-QNG for $\eta\leq 0.1$ indicates its ability to escape from local minima and to explore the widest region in the parameter space while searching for the global minimum.
To study the convergence behaviour of the optimization algorithms under consideration, in Fig. 1(b) and (d) we plot the mean (symbols) and the standard deviation (error bars) of the number of steps to convergence of the optimization procedure in a series of 200 trials, as a function of the stepsize value (learning rate, $\eta$). From Fig. 1(b) one can see that at $N=6$ Momentum-QNG demonstrates a decent convergence behavior over the whole range of stepsize values under consideration and even the best one for $\eta<0.075$. Again, on Fig. 1(d) for $N=12$ and $\eta>0.1$ Momentum-QNG shows a tendency to divergence, while for $\eta<0.075$ it demonstrates the best result. For further details of our calculations see [24].
@@
Figure 1: Benchmarking Momentum-QNG together with QNG, Momentum and Adam on the portfolio optimization problem. The vertical axis shows the mean (symbols) and the standard deviation (error bars) of the difference between the optimized and the ground state energy (a) ($N=6$), (c) ($N=12$) and the number of steps to convergence (b) ($N=6$), (d) ($N=12$) in a series of 200 trials, while the horizontal axis shows the stepsize value (learning rate, $\eta$) of four different optimizers under consideration. Each trial is initialized with random-guess values of the variational parameters, being the same for all optimizers. Then the optimization process runs for 200 steps (or until convergence) independently by each optimizer.

### v4

@@ Abstract
A Quantum Natural Gradient (QNG) algorithm for optimization of variational quantum circuits has been proposed recently. In this study, we employ the Langevin equation with a QNG stochastic force to demonstrate that its discrete-time solution gives a generalized form of the above-specified algorithm, which we call Momentum-QNG. Similar to other optimization algorithms with the momentum term, such as the Stochastic Gradient Descent with momentum, RMSProp with momentum and Adam, Momentum-QNG is more effective to escape local minima and plateaus in the variational parameter space and, therefore, demonstrates an improved performance compared to the basic QNG. In this paper we benchmark Momentum-QNG together with the basic QNG, Adam and Momentum optimizers and explore its convergence behaviour. Among the benchmarking problems studied, the best result is obtained for the quantum Sherrington-Kirkpatrick model in the strong spin glass regime. Our open-source code is available at https://github.com/borbysh/Momentum-QNG
@@ 1 Introduction
A computationally efficient method for evaluating analytic gradients on quantum hardware has been recently proposed [6]. Therefore, the application of optimization algorithms from the Stochastic Gradient Descent (SGD) family has become possible. However, the path of steepest descent in the parameter space, guided by the (opposite) gradient vector direction, is usually not optimal, because it depends on the number of variational parameters, which is usually excessive. The same overparametrization problem is present in classical Machine Learning (ML) [7]. To mitigate it, a Natural Gradient (NG) concept has been proposed [8]. Contrary to vanilla SGD, NG determines the steepest descent direction taking into account the Fisher Information Matrix, which consists of the components of the Riemannian metric tensor in the space of variational parameters. In this way, the optimization path becomes invariant under arbitrary reparametrization [8] and, therefore, does not suffer from overparametrization [9]. To speed up calculations, the Fisher Information Matrix is usually approximated with different methods (see e.g. Refs. [10, 11]).
Inspired by the NG approach, Stokes et al. [12] have recently proposed its generalization to the quantum circuit optimization task, which they called the Quantum Natural Gradient (QNG) optimizer. They considered a parametric family of unitary operators $U_{\theta}\in U(D)$, which are indexed by real parameters $\theta\in\mathbb{R}^{d}$. With a fixed reference unit vector $|0\rangle\in\mathbb{C}^{D}$ and a Hermitian operator $H=H^{\dagger}$ acting on $\mathbb{C}^{D}$, they consider the following optimization problem [12]:
@@
where $\psi_{\theta}=U_{\theta}|0\rangle$ and $P_{\psi_{\theta}}\in\mathbb{CP}^{D-1}$ is the associated projector. Note that $\psi_{\theta}$ is normalized, since $U_{\theta}$ is unitary. Any local optimum of the nonconvex objective function $\mathcal{L}(\theta)$ can be found by iterating the following discrete-time dynamical system [12],
$\theta_{n+1}=\operatorname*{arg\,min}_{\theta\in\mathbb{R}^{d}}\left[\langle\theta-\theta_{n},\nabla\mathcal{L}(\theta_{n})\rangle+\frac{1}{2\eta}\|\theta-\theta_{n}\|_{g(\theta_{n})}^{2}\right]\,$, (2)
where $\eta>0$ is a positively defined constant, $g(\theta_{n})$ is the symmetric matrix of the Fubini-Study metric tensor $g_{ij}(\theta)=\operatorname{Re}[G_{ij}(\theta)]$, and the Quantum Geometric Tensor is defined as follows (for further details see Stokes et al.[12] and references therein):
$G_{ij}(\theta)=\left\langle\frac{\partial\psi_{\theta}}{\partial\theta^{i}},\frac{\partial\psi_{\theta}}{\partial\theta^{j}}\right\rangle-\left\langle\frac{\partial\psi_{\theta}}{\partial\theta^{i}},\psi_{\theta}\right\rangle\left\langle\psi_{\theta},\frac{\partial\psi_{\theta}}{\partial\theta^{j}}\right\rangle\,+\lambda\delta_{ij},$ (3)
where the last term with the Kronecker delta $\delta_{ij}$ and $\lambda\geq 0$ is added for regularization purpose.
Note, that generally the Quantum Geometric tensor can be defined for more general parametrized normalized wave functions. In particular, it naturally appears in the stochastic reconfiguration method [13] to optimize general variational quantum states and was used in the optimization of Neural Quantum States [14, 15] and with certain modifications of some Tensor Network states [16]. Besides, such tensors appear in the imaginary time evolution of Gaussian and generalized Gaussian states [17].
In equation (2), according to Stokes et al.[12], the following notation is introduced:
$\|\theta-\theta_{n}\|_{g(\theta_{n})}^{2}=\langle\theta-\theta_{n},g(\theta_{n})(\theta-\theta_{n})\rangle\kern 5.0pt.$ (4)
Then, the first-order optimality condition corresponding to equation (2) is:
$g(\theta_{n})(\theta_{n+1}-\theta_{n})=-\eta\cdot\nabla\mathcal{L}(\theta_{n})\kern 5.0pt.$ (5)
@@
$\theta_{n+1}-\theta_{n}=-\eta\cdot g^{-1}({\theta_{n}})\nabla\mathcal{L}(\theta_{n})\kern 5.0pt.$ (6)
In this way, with the Fubini-Study metric tensor introduced, the implied descent direction in the parameter space, given by the right-hand side of equation (6), becomes invariant with respect to arbitrary reparametrization and, therefore, to the details of the quantum circuit architecture under consideration. With their QNG optimizer (6), the authors [12] achieved a considerable improvement in optimization performance compared to SGD and Adam [18].
However, being rather effective for convex optimization, QNG often sticks to local minima, saddles and plateaus of nonconvex loss functions. In classical ML applications, employing optimization algorithms with momentum (inertial) term, such as SGD with momentum [19], RMSProp with momentum [20] and Adam [18], has demonstrated better convergence characteristics.
Recently, Borysenko and Byshkin [21] demonstrated that a discrete-time solution of the Langevin equation with stochastic gradient force term results in the well-known SGD with momentum [19] optimization algorithm. In this paper, we study a particular case of Langevin dynamics with the QNG stochastic force term. In Section 2 we give a brief picture of the multidimensional discrete-time Langevin dynamics and show its relation to the stochastic optimization process. Based on these results, in Section 3 we derive a generalized QNG optimization algorithm, which we call Momentum-QNG and benchmark it together with the basic QNG, Momentum and Adam on several optimization tasks to demonstrate its improved performance.
2 Langevin dynamics and its relation to stochastic optimization
The adaptation of Langevin dynamics for optimization suggests a new prospective research direction [22, 23]. For optimization of analytically defined objective functions, even a quantum form of the Langevin dynamics has been proposed [24].
In Langevin dynamics, two forces are added to the classical Newton equation of motion: a viscous friction force proportional to the velocity with a friction coefficient $\gamma\geq 0$ and a thermal white noise. Explicitly, the Langevin dynamics of a (virtual) Brownian particle with unit mass $m=1$ in the space of real variables $\theta\in\mathbb{R}^{d}$ and real time $t$ can be described by the following equation (see e.g. Refs.[25, 26, 27, 28]):
@@ 2 Langevin dynamics and its relation to stochastic optimization
where $v=d\theta/dt$ denotes velocity, $f(\theta)\in\mathbb{R}^{d}$ is a regular force and $R(t)\in\mathbb{R}^{d}$ is a random uncorrelated force with zero mean $\left\langle R(t)\right\rangle=0$ and temperature-dependent magnitude:
$\left\langle R^{i}(t)\cdot R^{j}(t^{\prime})\right\rangle=2T\gamma\delta(t-t^{\prime})\delta_{ij},$ (8)
$\delta(t-t^{\prime})$ being the Dirac delta function.
Similar to the $d=3$ case, the temperature $T$ in the $d$-dimensional space of our variational parameters may be introduced as twice the mean kinetic energy $E_{k}$ of a particle divided by $d$:
$T=\frac{2E_{k}}{d}=\frac{1}{d}\sum_{i=1}^{d}\left\langle\left(v^{i}\right)^{2}\right\rangle.$ (9)
In equation (9) we set the Boltzmann’s constant $k_{B}=1$ for brevity.
The discrete-time form of equation (7) with stochastic force $\hat{f}=f+R$ reads:
@@
After changing to discrete time, equation (8) becomes:
$\left\langle R^{2}\right\rangle\Delta t=2T\gamma,$ (14)
where $\left\langle R^{2}\right\rangle=d^{-1}\cdot\sum_{i=1}^{d}\left\langle\left(R^{i}\right)^{2}\right\rangle$.
Using equation (12) to change variables from $\gamma$ to $\rho$, the thermostatic condition (14), corresponding to the discrete-time form of the Langevin equation (10), becomes:
$\left\langle R^{2}\right\rangle\Delta t^{2}=4T\cdot\frac{1-\rho}{1+\rho}.$ (15)
At the same time, in discrete variables, the temperature (9) becomes
$T=\frac{1}{d}\sum_{i=1}^{d}\left\langle\ \left(\frac{\Delta\theta^{i}}{\Delta t}\right)^{2}\right\rangle=\frac{1}{d\cdot\Delta t^{2}}\sum_{i=1}^{d}\left(\sigma^{i}\right)^{2}=\frac{\sigma^{2}}{\Delta t^{2}},$ (16)
where $\sigma=\sqrt{d^{-1}\cdot\sum_{i=1}^{d}\left(\sigma^{i}\right)^{2}}$ is a standard deviation magnitude in the space of variational parameters.
Now, with equations (13) and (16) in mind, one can derive from the thermostatic condition (15) the next relation:
$\sigma=\eta\cdot\sqrt{\frac{\left\langle R^{2}\right\rangle}{1-\rho^{2}}},$ (17)
which describes fluctuations of variational parameters in the vicinity of a loss function extremum under the influence of random noise. One can use equation (17) to estimate the accuracy of the optimized variational parameters.
From equation (17) one can see that the mean random jump length $\sigma$ increases as the momentum coefficient $\rho$ increases. Therefore, choosing $\rho>0$ helps to escape local minima and plateaus, where the local gradient values vanish.
3 Benchmarking of the quantum natural gradient descent with momentum
@@
from equation (11) one obtains a generalized form of the QNG optimizer (6):
@@
which we call Momentum-QNG. Note, that equation (19) reduces to the basic QNG (6) for $\rho=0$.
From equation (19) one can see that our Momentum-QNG optimization algorithm is a quantum adaptation of SGD with momentum [19] and should not be confused with the recently introduced Momentum QNG algorithm [29] (which is a quantum adaptation of Adam [18]) proposed by the authors of qBang [29].
In this section we benchmark four gradient-based optimization algorithms, integrated into the PennyLane [30] quantum computation package as Python3 classes, at different learning rate values with the rest hyperparameter values set as follows: Adam($\beta_{1}=0.9,\beta_{2}=0.99,\epsilon=10^{-8}$), Momentum($\rho=0.9$), QNG($\lambda=0.5$), Momentum-QNG($\rho=0.9,\lambda=0.5$). In QNG and Momentum-QNG, the Fubini-Study metric tensor (see equation (3)) is computed under the block-diagonal approximation and the regularization coefficient $\lambda=0.5$ is applied. In all calculations we set $\Delta\theta_{0}=0$ as the initial condition.
@@ 3.1.1 Investment Portfolio Optimization
Our first test-drive model is an Investment Portfolio Optimization task, which can be mapped to the $N$-particle (NN is a number of companies in portfolio) Ising spin-glass model. The ground state of the corresponding $N$-qubit Hamiltonian is parametrized with the VQE ansatz and further optimized to find the energy minimum. Here we explore the cases $N=6$, 11 and 12. To benchmark the performance of different optimizers, we run a series of 200 trials on a modified tutorial code by Chi-Chun Chen [31] for a range of learning rate values: $0.01\leq\eta\leq 3$. Each trial is initialized with random-guess values of variational parameters, being the same for all optimizers. Next, the optimization process runs for 200 steps or until energy convergence up to the 3-digit accuracy. As a result of each optimization run, we calculate $\Delta E=E_{\text{opt}}-E_{\text{ground}}$ – the difference between the optimized and the exact ground state energy.
To compare the performance of different optimizers, on Fig. 1(a), (c) and (e) we plot the mean (symbols) and standard deviation (shaded regions) values of $\Delta E$ as a function of the learning rate $\eta$. From Fig. 1(a), (c) and (e) one can see that for $N=6$, $N=11$ and $N=12$ all the three momentum-amended optimizers (Momentum-QNG, Momentum and Adam) give similar best results in their convergence domain and significantly outperform the momentumless QNG.
To study the convergence behaviour of the optimization algorithms under consideration, in Fig. 1(b), (d) and (f) we plot the mean (symbols) and the standard deviation (shaded regions) of the number of steps to convergence, as a function of the learning rate $\eta$. Again, for $N=6$, $N=11$ and $N=12$ all the momentum-amended optimizers demonstrate similar results in their convergence domain. The momentumless QNG shows the fastest convergence behaviour, though the highest energy misfit.
It is worth noting that Adam demonstrates the most robust performance with the widest convergence domain. For further details of our calculations see our code [32].
@@
Figure 1: Benchmarking Momentum-QNG together with QNG, Momentum and Adam on the portfolio optimization problem. The vertical axis shows the mean (symbols) and the standard deviation (shaded regions) of the difference between the optimized and the ground state energy (a) ($N=6$), (c) ($N=11$), (e) ($N=12$) and of the number of steps to convergence (b) ($N=6$), (d) ($N=11$), (f) ($N=12$) in a series of 200 trials, while the horizontal axis shows the learning rate $\eta$ of four different optimizers under consideration.
3.1.2 The Sherrington-Kirkpatrick model
In this section we optimize the quantum Sherrington-Kirkpatrick (SK) model [33] to find its ground-state energy in the framework of the VQE approximation. For a recent discussion of the quantum SK model and its ground state ansatzes see Ref.[34].
We consider the $N$-qubit quantum SK model in the transverse field, defined by the following Hamiltonian:
$\hat{H}=\sum_{i,j}J_{ij}\hat{\sigma}_{i}^{z}\hat{\sigma}_{j}^{z}-g\sum_{i}\hat{\sigma}_{i}^{x},$ (20)
where the first sum is taken over all pairs of sites $(ij)$ and $J_{ij}=\mathcal{N}(0,1)/\sqrt{N}$ are sampled from the normal distribution with zero mean and $1/N$ variance. At small $g<1.5$ the model is generally in the spin glass phase. At large $g$ the model becomes paramagnetic in the $x$ direction.
To illustrate the performance of different optimizers, we consider the VQE optimization problem aimed to minimize the ground state energy expectation value:
$E_{0}(\theta)=\langle 0|U^{{\dagger}}_{\theta}\hat{H}U_{\theta}|0\rangle,$ (21)
where $U_{\theta}$ is a parametrized unitary matrix.
In Fig. 2 below we demonstrate results for $N=8$ qubits. To benchmark the performance of different optimizers, we run a series of 200 trials based on our modified tutorial code [35] for a range of learning rate values: $10^{-3}\leq\eta\leq 10$. Each trial is initialized with random-guess values of variational parameters, being the same for all optimizers. Next, the optimization process runs for 300 steps or until energy convergence up to the 5-digit accuracy. As a result of each optimization run, we calculate the $\text{error}=100\%\cdot\left(E_{0}^{\text{true}}-E_{0}(\theta)\right)/E_{0}^{\text{true}}$ – the relative difference between the optimized and the true ground state energy.
Figure 2: Benchmarking Momentum-QNG together with QNG, Momentum and Adam on the Sherrington-Kirkpatrick model at three different values of transverse field $g$ (indicated at figure captions). The vertical axis shows the mean (symbols) and the standard deviation (shaded regions) of the difference (in percents) between the optimized and the true ground state energy (a) ($g=0.1$), (c) ($g=10^{-3}$), (e) ($g=10^{-5}$) and of the number of steps to convergence (b) ($g=0.1$), (d) ($g=10^{-3}$), (f) ($g=10^{-5}$) in a series of 200 trials, while the horizontal axis shows the learning rate $\eta$ of four different optimizers under consideration.
To compare the performance of different optimizers, on Fig. 2(a), (c) and (e) we plot the mean (symbols) and standard deviation (shaded regions) values of the ground energy error as a function of the learning rate $\eta$. From Fig. 2 one can see that for $g=0.1$ (a) the basic QNG achieves the least error value. For $g=10^{-3}$ (c) Momentum and Momentum-QNG demonstrate almost equal best results and for $g=10^{-5}$ (e) Momentum-QNG shows the best result. It is worth to note that Adam demonstrates modest optimization performance in all three cases.
To study the convergence behaviour of the optimization algorithms under consideration, in Fig. 2(b), (d) and (f) we plot the mean (symbols) and the standard deviation (shaded regions) of the number of steps to convergence, as a function of the learning rate $\eta$. One can see that the basic QNG demonstrates the fastest convergence in the most narrow domain, while the rest three optimizers behave similarly.
Our raw data and additional numeric results can be found at our project page [36].